OpenStudy (this):

The area of a rectangle is given by A = x2 + 4x – 5. Find the value of x if the total area of the rectangle is equal to 7 square units. a) x = 7 b) x = 5 c) x = 1 d) x = 2

terenzreignz (terenzreignz):

The best way to start these is to put down what we're given... What do we know so far?

terenzreignz (terenzreignz):

We know that the area of the rectangle is 7 right? What else do we know?

terenzreignz (terenzreignz):

Work with me here, you should be able to get to the answer by yourself...

OpenStudy (this):

ya im trying

terenzreignz (terenzreignz):

so, We know that A = 7 Therefore... we can substitute 7 for A in the first equation... what do we get?

terenzreignz (terenzreignz):

Ok, so we know that A = 7 right? Therefore.... 7 = x^2 + 4x - 5, do you get why?

OpenStudy (this):

o yes

terenzreignz (terenzreignz):

Can you solve Quadratic equations?

OpenStudy (this):

hun not really

terenzreignz (terenzreignz):

so we have 7 = x^2 + 4x - 5 What I want you to do is manipulate this equation so that there is a 0 on one side.

terenzreignz (terenzreignz):

It's actually simple :) All you have to do is add (-7) to both sides of the equation go ahead, try it, and show me what you got.

OpenStudy (this):

x^2+4x-12

OpenStudy (this):

im not sure just trying ;)

terenzreignz (terenzreignz):

yes, but that was not an equation, but close enough: x^2 + 4x - 12 = 0 Can you do it from here?

OpenStudy (this):

yes let me try

OpenStudy (this):